# Mitigating Omitted Variable Bias in Empirical Software Engineering

### v2

@@ Abstract.
This paper illustrates the impact of omitted variable bias on two case studies in the software engineering domain, and uses them to present methods to investigate the possible presence of omitted variable bias, to estimate its impact, and to mitigate its drawbacks. The analysis techniques we present are based on causal structural models of the variables of interest, which provide a practical, intuitive summary of the key relations among variables.
@@ 1. Introduction
Consider a common task of any empirical discipline: estimating the (average) effect that changing a variable $X$ has on another, dependent variable $Y$. In experimental study design parlance, $X$ is the treatment, and $Y$ is the outcome or response. Here is a concrete example in the domain of software engineering, which we will look into more closely in Section 3: estimating how using a different programming language (treatment $X$) affects the quality (outcome $Y$) of a program implemented in that language. If we have observational data about $X$ and $Y$, we may simply fit a statistical model—anything from a simple linear regression to fancier devices—on these data, and interpret the fitted model’s parameters connecting $X$ and $Y$ as an estimate of this effect. In this process, a well-known snag is omitted variable bias: if there exists another unmeasured variable $Z$ that affects both $X$ and $Y$, our estimate of the effect $X\to Y$ will spuriously include the combined effect of $X$ and $Z$ instead of the effect of $X$ alone. Continuing our empirical software engineering example, $Z$ may be a programmer’s intrinsic skills, which are all but sure to also affect the quality $Y$ of the written programs. In such a scenario, adding more data about $X$ and $Y$ (i.e., sampling more datapoints) is not going to help; in fact, it may just entrench our reliance on the biased estimate by reducing its variance and giving the false impression of reliability or “significance”.
@@
Contributions.
@@
In a nutshell, the main aims of this paper are methodological: the case studies of Section 3 and Section 4 are based on real data, but are not meant to lead to novel findings about their domains. Instead, they demonstrate plausible scenarios where confounding may occur, and to explore different mitigation strategies.
Organization.
@@
Section 2 first presents relevant related work on the topics of mitigating biases in statistical analysis and, more generally, best statistical practices for empirical software engineering; then, it introduces the key notations and concepts that the rest of the paper relies on. The rest of the paper demonstrates how to deal with omitted variable bias in two case studies: Section 3 uses a comparatively simpler model to estimate the effect of programming languages on code quality (also mentioned in the introduction), which serves as a relatively uncomplicated scenario. Then, Section 4 uses a more complex model to analyze the relation between team size (how many developers work on a project) and effort (how long it takes to complete the project). Although the two case studies differ in complexity, they are both based on realistic scenarios and models that we investigated in our previous work [14, 11]. Section 5 serves as a high-level summary of the whole article, presenting guidelines, in the form of a sequence of analysis steps that generate fundamental information to support the design of a study that mitigates omitted variable bias. Finally, Section 6 concludes the paper with a short summary of the main contributions.
2. Related Work and Background
@@
2.1. Related Work
One of science’s ultimate goals is understanding the processes that underlie observed phenomena. This means discovering cause/effect relations between variables, as opposed to mere statistical associations. While the roots of causal analysis date back to Neyman’s potential outcomes framework [45], a comprehensive understanding of causality has emerged only later, in the late part of the 20th century. The key milestones in the development of a robust understanding of causality include: (1) Rubin [40]built upon Neyman’s pioneering work, introducing a framework for causality in nonrandomized (observational) studies. (2) Angrist et al. [2]introduced instrumental variables [^1] In a nutshell, an instrumental variable is a variable that acts like a natural experiment on the treatment. based on the Neyman-Rubin framework. (3) Working at about the same time as Rubin, Pearl started focusing on structural (graph) models [35], which culminated in his celebrated techniques for rigorously analyzing causal effects and confounding Pearl [36].
Among Pearl’s work, directed acyclic graphs (DAGs) have become widely used to model the structural dependencies between observed (and unobserved) variables. As we will demonstrate already with the simple example of Section 2.2, DAGs are, first of all, a practical notation to specify causal relations. They also support techniques to estimate the strength of the causal relation among some nodes in the graph. Usually this is done by constructing a (linear) statistical model among variables, selected according to the DAG’s structure. (We will demonstrate this shortly in Section 2.3.) Nowadays, causal analysis based on DAGs is routinely used in disciplines with a strong empirical component such as medicine [54, 46], epidemiology [18], economics [23], and biology [24].
@@ 2.1. Related Work
In the last few years, some more empirical software engineering research was published, targeting varied topics such as: (1) modeling rules of human knowledge and how they are made available to artificial intelligence systems [21]; (2) studying the impact of social media posts on the popularity of open-source projects [10]; (3) analyzing dependencies in configurable software systems [20]; (4) studying the impact of programming languages on coding competitions [15].
@@
Figure 1. Three possible processes where variable $Y$ depends on variables $X$ and $Z$, and the corresponding DAGs capturing these structural relations.
@@
Figure 2. Two linear regression models that capture the dependence between $X$, $Y$, and $Z$.
2.2. Causal Dependencies and DAGs
@@
Process $p_{1}$ in 1(a) produces values of $X$ that are drawn randomly from a normal distribution with zero mean and unit standard deviation; and values of $Y$ that are perfectly proportional to those of $X$. Even in such an ideal scenario, any empirical measure of $X$ and $Y$ will include some measurement error $\epsilon$, which process $p_{1}$ models as another normal random variable that, together with $X$, determines the value of $Y$. We stress that we interpret 1(a)’s equations as capturing the causal dependencies among variables: $X$ and $\epsilon$ are drawn randomly (and independent of each other), and their random values determine (“cause”) the value of $Y$ in each draw. Correspondingly, the DAG (directed acyclic graph [36]) in 1(a) captures this causal relation between $X$ and $Y$ in a qualitative way: an edge connects $X$ to $Y$ to denote that the values of $X$ and $Y$ are related; furthermore, the edge is directed from $X$ to $Y$ to denote that changing $X$ directly affects $Y$, that is, it causes $Y$ to change.
Process $p_{2}$ in 1(b) involves a third variable $Z$. Just like variable $X$, $Z$’s values are drawn randomly from a normal distribution with zero mean and unit standard deviation. Then, the values of $Y$ are perfectly proportional to a linear combination of $X$ and ZZ—still with a term $\epsilon$ to account for measurement errors. Variables $X$ and $Z$ are independent of each other; the DAG in 1(b) clearly shows this independence, since it does not include any edge between $X$ and $Z$.
@@
2.3. Inference and Confounders
In an empirical study, the goal is to estimate, from a sample of the data, the parameters of a statistical model that captures the relations among observed variables. Let’s shift perspective on our illustrative examples: now, we are given a data sample $D_{1}$, $D_{2}$, $D_{3}$ respectively produced by each process $p_{1}$, $p_{2}$, $p_{3}$. Each data sample consists of many triples $(x_{i},y_{i},z_{i})$ of concrete values taken by $X$, $Y$, and ZZ [^2] Process $p_{1}$’s data sample only consists of pairs, since this process does not include a variable $Z$. in the $i$th observation—that is, the process’s $i$th draw. Now, we are imagining that we do not know the equations governing the generation process, but we want to quantitatively estimate the relation between $X$ and $Y$ from the observed sample.
@@ 2.3. Inference and Confounders
Table 1. Estimating the parameters of Figure 1’s processes $p_{1}$, $p_{2}$, and $p_{3}$ with Figure 2’s regression models $m_{1}$ and $m_{2}$.
@@
The case of process $p_{2}$ is more interesting, since we may analyze its data using either model $m_{1}$ or model $m_{2}$. As we expect, regression model $m_{2}$ accurately infers the value of parameter $\beta=b=0.40$, again reflecting the true dependency between $X$ and $Y$. Somewhat less obviously, the simpler model $m_{1}$ still infers the same correct value of parameter $\beta=b=0.40$, even if $Y$ also depends on $Z$ in generation process $p_{2}$. With model $m_{1}$, the effect of $Z$ on $Y$ has spilled into the estimate of the standard deviation $\sigma$ of the error term, which is in fact equal to $1.22$, greater than the “true” error standard deviation 1.01.0.333Equivalently, we can rewrite 1(b)’s generative equations as $Y=bX+E$, where $E\sim\mathrm{Normal}(0,\sqrt{1+c^{2}})$; if $c=0.7$, $\sqrt{1+c^{2}}\simeq 1.22$, which is exactly the inferred value of $\sigma$ in $m_{1}$ fitted on data from $p_{2}$.
@@
Using a model such as $m_{1}$ to analyze data from a process whose dependencies include a confounder (like process $p_{3}$) is an instance of omitted variable bias. The rest of the paper describes more realistic case studies where omitted variable bias may occur, and presents various mitigation strategies to counter the bias and recover a precise estimate of the causal effect linking treatment $X$ and outcome $Y$.
@@
2.4. Other Forms of Confounding
The term “confounding” is used to denote different, related concepts in the statistical analysis of empirical data [17].
  • In this paper, we use “confounding” to denote bias in the estimate of a causal effect—as we demonstrated in a nutshell in the previous sections. This notion of confounding is customary in modern causal analysis [36], and in related approaches to mitigate confounding, such as instrumental variables [2].
  • An early usage of “confounding” in statistics [49] was to denote noncollapsibility: roughly speaking, the association (correlation) between two variables is noncollapsible if its magnitude depends on whether we condition or not on another variable. Thus, noncollapsibility may denote a spurious causal effect [53], as in the previous meaning of “confounding”, but may also arise with different kinds of relations among variables; conversely, causal confounding may or may not determine noncollapsibility [17].
  • In classical frequentist statistics, the term “confounding” often denotes scenarios where main effects and interactions are inseparable [13]. If the interactions are not of interest, a study may deliberately introduce this kind of confounding, so that the main inferred effects summarize true effects and interactions.

### v3

@@ Abstract
This paper illustrates the impact of omitted variable bias on two illustrative examples in the software engineering domain, and uses them to present methods to investigate the possible presence of omitted variable bias, to estimate its impact, and to mitigate its drawbacks. The analysis techniques we present are based on causal structural models of the variables of interest, which provide a practical, intuitive summary of the key relations among variables.
@@ 1 Introduction
Consider a common task of any empirical discipline: estimating the (average) effect that changing a variable $X$ has on another, dependent variable $Y$. In experimental study design parlance, $X$ is the treatment [^1] The term treatment comes from the practice of randomized controlled clinical trials, where it strictly denotes a binary variable (treatment vs. control group) that is randomly assigned. In statistical parlance, however, the term is routinely applied more generally to indicate any variable whose effect on an outcome is being analyzed. Under this general meaning, a treatment variable may not be randomly assigned [34] (for example, in an observational study) or it may be continuous [32] instead of binary. In this paper, we use “treatment” with this relaxed, more general meaning—as a synonym of “exposure”. or exposure, and $Y$ is the outcome or response. Here is a concrete example in the domain of software engineering, which we will look into more closely in Section 3: estimating how using a different programming language (treatment $X$) affects the quality (outcome $Y$) of a program implemented in that language. If we have observational data about $X$ and $Y$, we may simply fit a statistical model—anything from a simple linear regression to fancier devices—on these data, and interpret the fitted model’s parameters connecting $X$ and $Y$ as an estimate of this effect. In this process, a well-known snag is omitted variable bias: if there exists another unmeasured variable $Z$ that affects both $X$ and $Y$, our estimate of the effect $X\to Y$ will spuriously include the combined effect of $X$ and $Z$ instead of the effect of $X$ alone. Continuing our empirical software engineering example, $Z$ may be a programmer’s intrinsic skills, which are all but sure to also affect the quality $Y$ of the written programs. In such a scenario, additional data about $X$ and $Y$ (i.e., sampling more datapoints) is not going to help; in fact, it may just entrench our reliance on the biased estimate by reducing its variance and giving the false impression of reliability or “significance”.
@@
As we discuss more broadly in Section 2.1, there has been a growing interest in adopting robust, modern statistical practices in empirical software engineering research—especially those based on causal analysis. This paper contributes to this line of research, focusing on the practical and widespread problem of omitted variable bias.
1.1 Contributions
@@
The main aims of this paper are methodological: the illustrative examples of Section 3 and Section 4 are based on real data, but are not meant to lead to novel findings about their domains. Instead, they demonstrate plausible scenarios where confounding may occur, and explore different mitigation strategies.
1.2 Scope and Limitations
This paper focuses on a specific, yet widespread, challenge in empirical software engineering: estimating and mitigating omitted variable bias when analyzing observational data. The techniques we illustrate use structural causal models (DAGs, or directed acyclic graphs) to model the possible underlying causal relations between variables; adjustment sets to determine which variables should be included in a statistical model to make its inferences consistent with the causal relations; and sensitivity analysis to systematically explore the robustness and validity of the inferred results relative to a range of causal assumptions. The methods we present are applicable both a posteriori—to analyze non-experimental data—and a priori—to inform the design of observational or quasi-experimental studies and to proactively outline the boundary of their validity.
These methods provide an accessible and practical framework for dealing with a common source of confounding in observational studies. However, they do not cover all confounding problems and are not suitable for every kind of empirical study design:
  • We focus on recovering causal effects, as opposed to optimizing predictive performance or model succinctness in a purely statistical setting; in related work, we demonstrated scenarios where these goals lead to different results [22].
  • The methods we present in the paper are well-suited for cross-sectional studies, where data from a population is collected at a specific point in time. These studies are very common in empirical software engineering. Section 2.5 outlines a few approaches that extend confounding analysis to time-dependent data, which arise in longitudinal or panel studies.
  • Omitted variable bias is only one of many different threats to the validity of empirical studies. Different kinds of confounding may require different statistical approaches [50]; more generally, other sources of bias (selection bias, measurement bias, …) may jeopardize a study’s design validity at different levels [82].
  • The DAG-based methods we consider in this paper have been gaining traction, as they support rigorous, yet intuitive foundations for causal modeling, and they are sufficiently general to accommodate a wide range of applications. However, other causal analysis frameworks exist—such as difference-in-differences, instrumental variables, regression discontinuity designs, matching methods, and synthetic control—that may be more suitable in certain contexts. Without going into details, a key difference of these approaches to causality is that they usually address confounding through preprocessing, design, or data structure (for example, relying on timing or instruments) rather than by modeling and adjusting within a statistical model as in DAG-based approaches [69]. A more detailed presentation of these alternatives is outside this paper’s scope, and is available elsewhere [38, 4].
  • The DAG-based methods we demonstrate in the paper rely on several technical assumptions including: the absence of causal loops (acyclicity of the DAG), probabilistic relations between variables (as opposed to purely deterministic relations, which can only be modeled indirectly [7]), and enough evidence to inform a plausible causal structural model. Section 2 presents the underlying methods in detail, compares them to other, related approaches, and outlines their relevance for empirical software engineering research.
1.3 Organization
@@
Section 2 first presents relevant related work on the topics of mitigating biases in statistical analysis and, more generally, best statistical practices for empirical software engineering; then, it introduces the key notations and concepts that the rest of the paper relies on. The rest of the paper demonstrates how to deal with omitted variable bias in two illustrative examples: Section 3 uses a comparatively simpler model to estimate the effect of programming languages on code quality (also mentioned in the introduction), which serves as a relatively uncomplicated scenario. Then, Section 4 uses a more complex model to analyze the relation between team size (how many developers work on a project) and effort (how long it takes to complete the project). Although the two illustrative examples differ in complexity, they are both based on realistic scenarios and models that we investigated in our previous work [21, 18]. Section 5 serves as a high-level summary of the whole article, presenting guidelines, in the form of a sequence of analysis steps that generate fundamental information to support the design of a study that mitigates omitted variable bias. Finally, Section 6 concludes the paper with a short summary of the main contributions.
2 Related Work and Background
@@
2.1 Background
One of science’s ultimate goals is understanding the processes that underlie observed phenomena. This means discovering cause/effect relations between variables, as opposed to mere statistical associations. While the roots of causal analysis date back to Neyman’s potential outcomes framework [68], a comprehensive understanding of causality has emerged only later, in the late part of the 20th century. The key milestones in the development of a robust understanding of causality include: 1. Rubin [63]built upon Neyman’s pioneering work, introducing a framework for causality in nonrandomized (observational) studies. 2. Angrist et al. [3]introduced instrumental variables [^2] In a nutshell, an instrumental variable is a variable that acts like a natural experiment on the treatment. based on the Neyman-Rubin framework. 3. Working at about the same time as Rubin, Pearl started focusing on structural (graph) models [57], which culminated in his celebrated techniques for rigorously analyzing causal effects and confounding Pearl [58].
Among Pearl’s work, directed acyclic graphs (DAGs) have become widely used to model the structural dependencies between observed (and unobserved) variables. As we will demonstrate already with the simple example of Section 2.2, DAGs are, first of all, a practical notation to specify causal relations. [^3] An alternative approach to modeling causal relations is structural equation modeling (SEM). Semantically, Pearl’s causal DAG framework and SEM with latent variables are closely related. Specifically, Pearl’s original work built on SEM, and modern presentations of the causal DAG framework explicitly connect it to SEM [8]. Methodologically, however, the two approaches differ in their focus: causal DAGs (especially as we use them in the paper) specialize in modeling structural causal relations among (observed or latent) variables; SEM, instead, tends to mix measurement aspects (e.g., the relation between latent constructs and indicators) and structural ones (e.g., the relation between different constructs). They also support techniques to estimate the strength of the causal relation among some nodes in the graph. Usually this is done by constructing a (linear) statistical model among variables, selected according to the DAG’s structure. (We will demonstrate this shortly in Section 2.3.) Nowadays, causal analysis based on DAGs is routinely used in disciplines with a strong empirical component such as medicine [81, 70], epidemiology [27], economics [37], and biology [43].
@@ 2.1 Background
In the last few years, some more empirical software engineering research was published, targeting varied topics such as: 1. modeling rules of human knowledge and how they are made available to artificial intelligence systems [31]; 2. studying the impact of social media posts on the popularity of open-source projects [17]; 3. analyzing dependencies in configurable software systems [29]; 4. studying the impact of programming languages on coding competitions [22].
More broadly, these techniques belong to a broad and growing body of work on modernizing statistical practices [50]. This work has been the response to the realization that some “traditional” and widely used statistical methods have limitations in terms of interpretability and real-world significance [79, 1, 53]. While improving statistical practices is only a piece of the puzzle, inadequate methods have certainly concurred to the issues with replicability experienced in empirical research areas as diverse as medicine [39], psychology [14], and economics [10]. Since empirical software engineering research shares challenges and practices with these domains, adopting more robust data analysis practices will benefit the quality and impact of our research field [25, 42, 48, 55, 41].
@@
Figure 1: Three possible processes where variable $Y$ depends on variables $X$ and $Z$, and the corresponding DAGs capturing these structural relations.
@@
Figure 2: Two linear regression models that capture the dependence between $X$, $Y$, and $Z$.
2.2 Causal Dependencies and DAGs
@@
Process $p_{1}$ in 1(a) produces values of $X$ that are drawn randomly from a normal distribution with zero mean and unit standard deviation; and values of $Y$ that are linearly proportional to those of $X$. Even in such an ideal scenario, any empirical measure of $X$ and $Y$ will include some measurement error $\epsilon$, which process $p_{1}$ models as another normal random variable that, together with $X$, determines the value of $Y$. We stress that we interpret 1(a)’s equations as capturing the causal dependencies among variables: $X$ and $\epsilon$ are drawn randomly (and independent of each other), and their random values determine (“cause”) the value of $Y$ in each draw. Correspondingly, the DAG (directed acyclic graph [58]) in 1(a) captures this causal relation between $X$ and $Y$ in a qualitative way: an edge connects $X$ to $Y$ to denote that the values of $X$ and $Y$ are related; furthermore, the edge is directed from $X$ to $Y$ to denote that changing $X$ directly affects $Y$, that is, it causes $Y$ to change.
Process $p_{2}$ in 1(b) involves a third variable $Z$. Just like variable $X$, $Z$’s values are drawn randomly from a normal distribution with zero mean and unit standard deviation. Then, the values of $Y$ are a linear combination of $X$ and ZZ—still with a term $\epsilon$ to account for measurement errors. Variables $X$ and $Z$ are independent of each other; the DAG in 1(b) clearly shows this independence, since it does not include any edge between $X$ and $Z$.
@@
2.3 Inference and Confounders
Consider an empirical study whose goal is to estimate, from a sample of the data, the parameters of a statistical model that captures the relations among observed variables. Let’s shift perspective on our illustrative examples: now, we are given a data sample $D_{1}$, $D_{2}$, $D_{3}$ respectively produced by each process $p_{1}$, $p_{2}$, $p_{3}$. Each data sample consists of many triples $(x_{i},y_{i},z_{i})$ of concrete values taken by $X$, $Y$, and ZZ [^4] Process $p_{1}$’s data sample only consists of pairs, since this process does not include a variable $Z$. in the $i$th observation—that is, the process’s $i$th draw. Now, let’s assume that we do not know the equations governing the generation process, but we want to quantitatively estimate the relation between $X$ and $Y$ from the observed sample.
@@ 2.3 Inference and Confounders
Table 1: Estimating the parameters of Figure 1’s processes $p_{1}$, $p_{2}$, and $p_{3}$ with Figure 2’s regression models $m_{1}$ and $m_{2}$.
@@
The case of process $p_{2}$ is more interesting, since we may analyze its data using either model $m_{1}$ or model $m_{2}$. As we expect, regression model $m_{2}$ accurately infers the value of parameter $\beta=b=0.40$, again reflecting the true dependency between $X$ and $Y$. Somewhat less obviously, the simpler model $m_{1}$ still infers the same correct value of parameter $\beta=b=0.40$, even if $Y$ also depends on $Z$ in generation process $p_{2}$. With model $m_{1}$, the effect of $Z$ on $Y$ has spilled into the estimate of the standard deviation $\sigma$ of the error term, which is in fact equal to $1.22$, greater than the “true” error standard deviation 1.01.0.555Equivalently, we can rewrite 1(b)’s generative equations as $Y=bX+E$, where $E\sim\mathrm{Normal}(0,\sqrt{1+c^{2}})$; if $c=0.7$, $\sqrt{1+c^{2}}\simeq 1.22$, which is exactly the inferred value of $\sigma$ in $m_{1}$ fitted on data from $p_{2}$.
@@
Using a model such as $m_{1}$ to analyze data from a process whose dependencies include a confounder (like process $p_{3}$) is an instance of omitted variable bias. The rest of the paper describes more realistic illustrative examples where omitted variable bias may occur, and presents various mitigation strategies to counter the bias and recover a precise estimate of the causal effect linking treatment $X$ and outcome $Y$.
@@
Correction and sensitivity analysis.
Broadly speaking, the rest of the paper illustrates two complementary approaches to deal with omitted variable bias when analyzing observational data. First, we will present techniques that can correct for possible omitted variable bias. Based on structural models of causality (see Figure 1’s DAGs), such techniques suggest how to model the data in a way that filters out confounding effects and recovers more precisely the strength of actual causal relationships.
As we will see in our examples, correcting for possible confounders is not always possible or practical. On the one hand, some confounding variables may simply be inaccessible; on the other hand, the model of causality that underlies the application of adjustment techniques may itself be imprecise or uncertain. In such cases, a sensitivity analysis can mitigate uncertainty and clarify the boundary of validity of our findings. In a nutshell, a sensitivity analysis is a sort of “what if” analysis that quantifies the robustness of the results under different assumptions about the underlying process. [^6] It is important to notice that, despite their superficial similarities, the practice of sensitivity analysis and the malpractice of data dredging (also known as “$p$-hacking”) have opposite goals: “In $p$-hacking, many justifiable analyses are tried, and the one that attains statistical significance is reported. In sensitivity analysis, many justifiable analyses are tried, and all of them are described.” [50, p. 319]. A sensitivity analysis can corroborate results and expose limits of validity, but is not expected to give any kind of definitive, binary answer.
2.4 Other Forms of Confounding
The term “confounding” is used to denote different, related concepts in the statistical analysis of empirical data [26]. In this paper, we use “confounding” to denote bias in the estimate of a causal effect—as we demonstrated in a nutshell in the previous sections. This notion of confounding is customary in modern causal analysis [58], and in related approaches to mitigate confounding, such as instrumental variables [3].
An early usage of “confounding” in statistics was to denote noncollapsibility [26, 73]. In a nutshell, the effect of $X$ over $Y$ is collapsible over $Z$ if the marginal association between $X$ and $Y$ (obtained by averaging over $Z$) is the same as their conditional association (obtained by conditioning on $Z$); noncollapsibility holds when the two associations differ. While noncollapsibility is a purely statistical notion, the notion of confounding is rooted in causal relations, which depend not only on data but also on the data generation process [58]; therefore, the two concepts overlap but are distinct. In particular, noncollapsibility may signal causal confounding (for example, in a scenario like 1(c)); but it may also arise from nonlinearities without confounding. For example, consider a scenario with the same causal structure $X\to Y\leftarrow Z$ as in 1(b) but where $Y$ has a non-linear logistic relation with the linear combination of $X$ and $Z$. Since the causal structure is the same as in 1(b) there is no confounding; however, the effect of $X$ on $Y$ depends on whether we condition on $Z$, and hence there is noncollapsibility.
In classical frequentist statistics, the term “confounding” denotes an experimental design that makes two effects indistinguishable from data [20]. This usage of the term refers to statistical identifiability, and does not have, in general, any relation to causal effects. For example, a so-called fractional factorial design omits some combinations of factor levels, which may introduce confounding in this sense.